OpenStudy (anonymous):

how do i find out the vertex of a parabola knowing just this: y= -5x^2 ?

OpenStudy (anonymous):

(0,0)

OpenStudy (anonymous):

a = 5 b = 0 x = -b/a x = 0 y = -5(0)^2 y = 0 (0,0)

OpenStudy (anonymous):

I mean x = -b/(2a)

OpenStudy (anonymous):

i’m sorry what i didn’t understand that

OpenStudy (anonymous):

Do you know there is a formula for the vertex of a parabola?

OpenStudy (anonymous):

really? what is it

OpenStudy (anonymous):

http://www.algebra-class.com/vertex-formula.html x = -b/(2a), where the equation is ax^2 + bx + c

OpenStudy (anonymous):

i mean 2nd function

OpenStudy (anonymous):

Have you learnt derivative?

OpenStudy (anonymous):

no i haven’t

OpenStudy (anonymous):

ok for this one, it's obvious that the vertex is at (0,0). For other quadratic equation, use complete the square

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can you just give me a step-by-step tutorial on how to find out the vertex? maybe if it has something to do with the mirror axis or w.e.

OpenStudy (anonymous):

Or you mean the xavier method, right?

OpenStudy (anonymous):

\[5x ^{2} = (5x)(x) \]

OpenStudy (anonymous):

\[(5x)(x) = 0\] x = 0

OpenStudy (anonymous):

idk what xavier method is

OpenStudy (anonymous):

just tell me verbally what i need to do plz.

OpenStudy (anonymous):

First you find their x-intercepts Then add them together, divided by 2. You get the axis of symmetry. Plug the x value back to the original equation and you get y Then, the vertex is (x, y)

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

you are a very nice person

OpenStudy (anonymous):

wait but what do you mean “add Them” TOGETHER? what is tHEM???

OpenStudy (anonymous):

ohh the 2 intercepts nvm sorry

OpenStudy (anonymous):

but what if 

OpenStudy (anonymous):

it only has 1 intercept

OpenStudy (anonymous):

how do i know if i should add it together

OpenStudy (anonymous):

Then there's only 1 intercept. For example for this one, zero + zero, then divided by 2 it's still zero

OpenStudy (anonymous):

i know but..

OpenStudy (anonymous):

look

OpenStudy (anonymous):